OpenStudy (anonymous):

Evaluate 6(x+3) =2 (2x+5)

OpenStudy (anonymous):

Okay, lets first start out by distributing.

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

iHeartYou, we are not supposed to just post answers.

OpenStudy (anonymous):

@GOODMAN sorry

OpenStudy (anonymous):

Thats alrite, just dont do it again :D

OpenStudy (anonymous):

Trying to get a link eh @GOODMAN

OpenStudy (anonymous):

A link?

OpenStudy (anonymous):

Forget it

OpenStudy (anonymous):

Evaluate this nah a to the power of 5 x by a to the power of -5

OpenStudy (anonymous):

Here: 6(x+3) =2 (2x+5) 6x+18=4x+10

OpenStudy (anonymous):

I didnt understand..

OpenStudy (anonymous):

Can you please paraphrase?

OpenStudy (studyhard):

6x+18=4x+10 right? Okay, from here. Place all like terms together. 6x-4x=10-18 2x=-8 x=-4 That's how you do it! :) hope you get it